OpenStudy (anonymous):

A rock is dropped from a cliff and hits the ground 6.0 seconds later. How high is the cliff?

OpenStudy (shamim):

Use h=ut+.5gt^2

OpenStudy (michele_laino):

Hint: the relationship between the height and the time, is: \[h = \frac{1}{2}g{t^2}\] where g is the gravity, namely g=9.81 m/sec^2. So substituting your data into the equation above, we get: \[h = \frac{1}{2}g{t^2} = \frac{1}{2} \times 9.81 \times 36 = ...?\]